OpenStudy (anonymous):

David spends 1/2 of an hour on his math homework. This is 1/4 of the total ti David spends on his homework. Which expression can help find the total minutes David spends doing homework? a. (1/2 / 1/4) / 60 b. (1/2 / 1/4) x 60 c. (1/4 / 1/2) / 60 d. (1/4 / 1/2) x 60

OpenStudy (anonymous):

1/2 of an hour to do math, and that's 1/4 of the total time of homework. that means total time of homework is 4 times greater than 1/2 of an hour. 1/2 * 4 (hours)

OpenStudy (anonymous):

David spends = (1/2)*60 = 30 minutes.. Let David spends total time on homework is : x minutes Then, (1/4)*x = 30 x/4 = 30 x = 120 minutes.. Let us check the Choice B.. \[\frac{\frac{1}{2}}{\frac{1}{4}} \times 60 = \frac{4}{2} \times 60 = 2 \times 60 = 120\] So, did you get what will be the answer???

OpenStudy (anonymous):

I'm trying to figure it out, but I just don't get it

OpenStudy (anonymous):

Where is the problem you are facing??

OpenStudy (anonymous):

Why am I multiplying 60 and not dividing?

OpenStudy (anonymous):

I hour = 60 minutes.. So, 1/2 of an hours means: \[(1/2) \times 1Hour\] And 1 Hour = 60 minutes So, \[(1/2) \times 60 minutes..\] I can solve it by one more method can I show you???

OpenStudy (anonymous):

yes, please.

OpenStudy (anonymous):

Thank you for taking the time to help and explain this to me.

OpenStudy (anonymous):

David spends = 1/2 Hours... And this 1/2 is 1/4 of the total time he spends.. Let the total time be: x So, 1/2 = (1/4)*x = x/4 Multiply by 4 both the sides, 2 = x So, x = 2 NOTE: this is in hours.. So, total time spends on homework = 2 Hours As 1 Hour = 60 minutes So 2 hours = 60*2 = 120 minutes.. See, you have noted or not but I tell you one thing, READ THE QUESTION FIRST MORE CAREFULLY AND ENTIRELY.. In question, it is to be found FIND THE TOTAL MINUTES DAVID SPENDS DOING HOMEWORK.. So, then answer will be in MINUTES and not in HOURS.. Please note this point...

OpenStudy (anonymous):

I got it now. Thank you so much for you help!!!

OpenStudy (anonymous):

Welcome dear...